UT 07 Sep 2021 - Night #6 of Restart



Night Staff:

OSA: S. Allanson

ISp: E. Solheid

ISA: O. Kuhn

Technical Wizard: J. Hill (Tucson)

Mtn Manager: M. Wagner

Sunset: 18:42

Twlight Ends: 20:01

Twilght Starts: 04:34

Sunrise: 05:54

Starting Config: LBCs

Night Log:

Summary:
  • LBCB+LBCR:
    • flats taken at Us,B,V (LBCB) + R,I (LBCR)
    • attempted phot field, but bias levels were sometimes ~0 (see below)
  • MODS:
    • took data for rotator center measurement
    • took data to measure RZP. Logs are in /home/lbto/data/20210907/{l,r}dg/
      • updated left RZP and obtained transform data on 19 stars, mostly in good ~1" seeing but it was puffing up to 1.3-ish near the end. transform_collect wanted to take MODS1 red and MODS2 blue images, but MODS1 red headers are missing info (related to IT 8466). GCS & PCS info (but not the MODS version is there so hopefully that is enough
    • imaging dither test
    • long slit acq test
  • Critical IT 8465: LBCR bias/background levels drop to 0 for some images
  • Critical IT 8468: MODS2 mask error (grabber disengaged from mask?)
  • modsAlign on robs - ran into an issue when trying to edit stars.

01:13 Turned on the LBC. Taking 2 test biases.

01:15 Running up MODS. mods1r has filename with index number of only 3 digits, not 4. Also MODS1: m1.ie No route to host. mods1 status looked fine - but something must be down, looking... It looks like the IE were shut off during the day (perhaps related to pumping?). When the inst electronics are shut down, the IE service has to be cycled. Did that and it is looking good now.

01:29 Running the ccdInit.pro scripts for MODS1 & MODS2 - these set up the CCD controller state - usually we don't do this, but at restart good. It does not help the index number 3 digits problem. We had the 3-digit index number problem before, June 2020. Jerry found that M1.RC was not running the official IC program but instead a development version.

02:01 Starting Us+F flats. Us has 17k counts, F only 6000. Scale factor is 1:2.2 but I'll have to change this to 1:6.5

02:05 Starting Us+F at PA=180 and updated scale.

02:08 Test B + R at PA=180. Running series of flats with 5x scale. Counts are OK --- 22k-12k (B) and 33k-22k (R) --- but lots of out-of-focus stars.

02:15 Test V + I at PA=180. Running series with 10x scale. 23k-11k(V) and 33k-16k (I). Rotate and scale by 15x.


02:30 Uploading a DSS image in the TPE on robs is too slow due to the ongoing network issues - will work on OT on laptop & transfer files.


02:33 Elliott is going up to top-off MODS.

03:00 Done

LBC Photometric Standards

03:00 Clem+Landolt J175954+452015 field UBVRI

03:01 dohybrid. Seeing is very good, ~0.6" (est. seeing) - the blue converged in 4 iterations but the red is bouncing around a bit in z4. face-back ~ -0.1-0.2

03:13 copointing: b 031446 and r 031441
COPOINTING: B=31446 R=31441<br />Pointing updates: delta_IE = -1.23", delta_CA = 10.38"<br /> Mirror updates:  dX(mm)  dY(mm)  dRX(")  dRY(")<br />             SX:   0.52    0.22    4.63  -11.16<br />             DX:  -0.53   -0.26   -5.57   11.44<br />----------------------------------------------------------------------

03:18 Starting observation to take a series of 5 dithers and 2 images/dither at each of the filters UBV(LBCB) + VRI(LBCR). This should take about 30 minutes.

03:30 Meanwhile, running the simSnap on both MODS. The temps are all within the user interface alert thresholds (no error on the HK page).

03:40 A few LBCR images have 0 for the background. 033818 for example, but the next one, 033900 is OK. 033953 once again has zeros. Biases from the start of the night look fine, ~300 counts, but now the LBCR bias has 0 (see lbcr 034619). This behavior throws a question mark on our photometric standard star observations... Around this time, I also noted that the collimation was falling apart; on LBCR, the stars were donuts but it was not as bad on LBCB.

03:47 Turning off the LBCs... and back on.

03:51 TMS - Heejoo and John have been trying to get TMS started, but the MultiLine software is not coming up.

03:53 Taking 2 test biases with the LBCs. The first LBCR biases have noise. Taking a series of 20 - these are looking better. Nevertheless, the behavior is odd and should be investigated (critical IT 8465)

04:00 Reconfiguring for MODS. Noticed that MODS1-R exposure sieveSnap exposure had hung. That was my error... I had not restarted the IC program on M1.RC after quitting to take a screenshot to send to Jerry re: the 3-digit index numbers and a possible development vs official IC version (IT 8466).

04:45 Pointing and collimating near M71. While the guider was in acquire mode, we saw the left side drifting... but maybe the probe had not homed correctly though that doesn't explain why stars were drifting. Re-homed probes and we're pointing & collimating.

MODS center of rotation on detector

04:56 execBinoMODS M71rotcen_active0.txt followed by a series of track presets at 0, 180, 90 and finally 270.
  M1B M1R M2B M2R
active PA=0 5 3 7 7
track PA=0 6 4 8 8
track PA=180 7 5 9 9
track PA=90 8 6 10 10
track PA=270 9 7 11 11
05:18 There is some cirrus

MODS rotator zero points

Gaia 20p25 and Stone O

05:20 acqBinoMODS 20p25_acq.txt --- in the WFS images, the pupil is down at the bottom of the subimage, especially for the left side (look for images around UT 05:24). John requests that we do some on-axis collimation sometime tonight.

05:26 header is corrupted on MODS1R images - modsAlign cannot determine which has the slitmask in and which has it out. The headers are missing information about not only the mask position, but also the environment, instrument configuration - a lot is missing. I edited the MASKPOS and SLITMASK but still modsAlign complains. We'll acquire MODS1 masks in the blue.

05:41 modsAlign gives an error when trying to add a star (this was on robs). On obs4, it works (offset was offsetpointing -0.1011 4.085 -1.688 detxy rel)

Traceback (most recent call last):
File "/lbt/lbto/mods/bin/modsAlign", line 2962, in <module>
xNew,yNew = editStars(fieldImage,disp,xStar,yStar,verbose=verbose)
File "/lbt/lbto/mods/bin/modsAlign", line 1523, in editStars
elif cursKey == KeyMarkDelete:
NameError: name 'KeyMarkDelete' is not defined

05:46 Repeat acquisition, this time doing a Blue Acq for MODS1. MODS2 hung on Red partner command (05:50).

MODS1: modsAlign Masks/mods.3.g2025ep2.mms mods1b.20210907.0011.fits mods1b.20210907.0010.fitsmods1: offsetpointing -0.1574 -0.343 0.353 detxy rel

MODS2: modsAlign Masks/mods.3.g2025ep2.mms mods2r.20210907.0014.fits mods2r.20210907.0015.fits offsetpointing -0.1022 4.181 -1.691 detxy rel

06:02 - Confirmatory thru-slit images: mods1b 12 and mods2r 16 are ok, although on mods2r they are a bit to the lower right near the top and left sides and to the upper right in bottom right quadrant. Compared to the average rotational offset from all MOS alignments in 2020-2021 (Com20210531): -0.17(0.08) for MODS1 and -0.03(0.07) for MODS2, MODS1 is consistent but MODS2 is questionable.

06:10 acqBinoMODS StoneO _acqblue.txt StoneO _acq.txt

Now, I have a modsAlign error on obs4! I tried the same command on robs and got the same error. I repeated modsAlign on obs4 and rejected the box at the bottom left which has two equally bright stars in it - this worked.

on obs4:

MODS1B field image mods1b.20210907.0014.fits
Traceback (most recent call last):
File "/home/modseng/bin/modsAlign", line 2883, in <module>
xPeak,yPeak,maxPix = findPeak(fieldImage,xPix,yPix,searchRad)
File "/home/modseng/bin/modsAlign", line 736, in findPeak
maxPix = np.amax(image[ymin:ymax,xmin:xmax])
File "/lbt/mods_runtime/anaconda/lib/python2.7/site-packages/numpy/core/fromnumeric.py", line 2140, in amax
out=out, keepdims=keepdims)
File "/lbt/mods_runtime/anaconda/lib/python2.7/site-packages/numpy/core/_methods.py", line 26, in _amax
return umr_maximum(a, axis, None, out, keepdims)
ValueError: zero-size array to reduction operation maximum which has no identity
Exit 1

on robs:

MODS1B field image mods1b.20210907.0014.fits
Traceback (most recent call last):
File "/lbt/lbto/mods/bin/modsAlign", line 2884, in <module>
xPeak,yPeak,maxPix = findPeak(fieldImage,xPix,yPix,searchRad)
File "/lbt/lbto/mods/bin/modsAlign", line 741, in findPeak
maxPix = np.amax(image[ymin:ymax,xmin:xmax])
File "/usr/lib64/python3.7/site-packages/numpy/core/fromnumeric.py", line 2505, in amax
initial=initial)
File "/usr/lib64/python3.7/site-packages/numpy/core/fromnumeric.py", line 86, in _wrapreduction
return ufunc.reduce(obj, axis, dtype, out, **passkwargs)
ValueError: zero-size array to reduction operation maximum which has no identity

MODS1: modsAlign Masks/mods.36.StoneO.mms mods1b.20210907.0013.fits mods1b.20210907.0014.fits: offsetpointing -0.1282 -0.358 0.219 detxy rel

MODS2: modsAlign Masks/mods.36.StoneO.mms mods2r.20210907.0017.fits mods2r.20210907.0018.fits: offsetpointing -0.0307 4.499 -1.517 detxy rel

Confirmatory thru-slit images are 15 (M1) and 19 (M2) look ok - M2 stars bit low in top boxes, ok in bottom...

06:35 acqBinoMODS 00p25_acqblue.txt 00p25_acq.txt (on this field, the pupil is near the top of the subimage (guide probe Y -178 - before Y was ~ -150 or so, not extreme).

MODS1: (collimation is very poor): modsAlign Masks/mods.2.g00d25ep.mms mods1b.20210907.0016.fits mods1b.20210907.0017.fits: offsetpointing -0.1617 -0.602 0.399 detxy rel

MODS2: modsAlign Masks/mods.2.g00d25ep.mms 210903M8.06j.fits mods2r.20210907.0021.fits: offsetpointing -0.0170 4.561 -1.648 detxy rel

Confirmatory thru-slit images M1 18 and M2 22 look good... M2 still shows some sign of a scale error as the stars are a bit low in the top boxes and centered in Y in the bottom ones.

Average rotational offset for the 3 MOS alignments tonight:

MODS1 -0.1617,-0.1282, -0.1574 -> -0.1491 ~ -0.15 (will make an update to PCSInstrument.conf)

MODS2: -0.0170, -0.0307, -0.1022 -> -0.0500 (no change to be made)

as tcs@tcs1 in /lbt/data/config/tcs/PCS:

cp PCSInstrument.conf PCSInstrument.conf.20210907

edit PCSInstrument.conf to change MODS LEFT ROTATOR ZEROPOINT from 27.63 to 27.78.

Spectrophotometric Standard, BD+28 4211

07:00 acqBinoMODS bd284211_blue.acq bd284211.acq

MODS2: b19 -> offsetxy -1.663 11.835 rel --> b20, star is on left edge, 1.018 offset given, b21

MODS2: r23 -> offsetxy 3.654 7.982 rel --> r24, pretty good, gave -0.28" offset --> r25

07:14 execBinoMODS bd284211.obs
  M1B M1R M2B M2R
dual grating 22-24 10-12 12-14 26-28
red grating   13-15   29-31
blue grating 25-27   15-17  
dual prism 28-30 16-18 18-20 32-34
red prism   19-21   35-37
blue prism 31-33   21-23  
The counts are near or at saturation for MODS2B dual - seeing is very good, 0.53" on the guider

SX/DX seeing: 0.78"/0.53" but on the WFS, 0.54/0.53". Pupil is near the bottom of the subimage on both. gprobe at X = -70, Y = -121

Back to the RZP and transform

08:04 Bouncing PCS after updating PCSInstrument.conf (left rzp from 27.63 to 27.78)

Confirming the change was correct:

08:05 acqBinoMODS 00p25_acqBlue.txt 00p25_acq.txt

MODS1: 34 & 35 --> offsetpointing -0.0018 0.400 0.156 detxy rel --> 36, good

MODS2: 38 & 39 -> offsetpointing -0.0223 4.333 -1.594 detxy rel --> 40, good (still scale ?)

08:25 acqBinoMODS StoneO _acqBlue.txt StoneO _acq.txt

MODS1: 37 & 38 --> offsetpointing -0.0191 0.129 -0.206 detxy rel --> 39 looks pretty good

MODS2: 40 & 41 --> offsetpointing 0.0146 4.443 -1.634 detxy rel --> 42

Transform_collect

We should try to use MODS1 Blue since MODS1 Red has missing header information. Made a blue-only version of transformInit_PointingMask.pro and ran it on MODS1. MODS2 uses dual imaging.

Sending a binocular preset from IDL to Stone_A1, TS=15, GS=0, PA=180.

08:50 SX primary mirror panicked. 08:57 Steve recovered it & trying again, but no star found. Steve is checking the pointing. He'll also collimate on-axis so we can note the pupil position on the WFS subimage.

09:01 Collimating on-axis from 09:03 - 09:04 --- the pupil looks pretty well centered in the subimage cutout.

09:04 Trying the preset again - failed. Trying an on-axis preset and that worked. Now trying gs=1 and ts=15 for PA=180. That worked - gs=0 may have been just at the lower limit of the guide probe.

09:11 Taking test exposures of 15-sec. IMCS jumps because IMCS was locking as we took the exposure (expected - not error). Try 9-sec: m1b 41 (poor IQ), 42 (15-sec) and 43 (9-sec) better and m2b,r = 25,45

Use 9-sec:

09:20 Starting transform data collection. transform_collect, exp=9.0 (Stone_A1_star15_pa180.list - 8 stars) on both sides. Seeing has gone off a bit, ~1-1.2".

Transform_collect wants to use the red channel on MODS1 and the blue channel on MODS2. Sending MODS1 into dual-channel mode and starting the script anew on MODS1. MODS1R headers contain GCS & PCS guide star data, like GSMEAS_X,Y but not Acquisition Guide and WFS System Configuration, like AGW{X,Y}GP
  M1R M2B
     
1 24 28
2 25 29
3 26 30
4 27 31
7 28 32
8 29 33
9 30 34
16 31 35 - fuzzy extended source in guider
on-axis 32 36
1   37, 38 left is behind, so restarting
2   39
3   40
4   42
7   43, stopped with cntl-c after this
     
     
     

Trying to load a new list for TS=6 (6 stars). Problems sending preset - first, it does not like sending a bino preset from the session used for transform collect ,even though iif_register, /tcs was done. Starting the left side from another IDL session worked
  M1R M2B Comments
       
0 err err restarting, skipping star 0
1 37 49 seeing >~1.2-1.3"
2 38 50  
8 39 51  
11 40 52  
12 41 53  
13 42 54  
on-axis 43 55  

New list, Stone_A1_star28_pa180.list (5 stars).
  M1R M2B  
on-axis 44 56  
BG 45 57  
7 46 58  
15 47 59  
22 48 60  
29 49 61  
30 50 62 seeing is ~1.3, even 1.5"
on-axis 51 63  
Imaging with dithers

11:23 execBinoMODS M76_abbrev.txt

This script takes a set of 6 dithered images at : (0,0), (35,35), (-35,35), (-35,-35), (35,-35), (0,0) using aysnc offsets (sync offsets are discouraged in imaging scripts).

Gave an error on the left side and a warning on the right (error > 50" and warning: 40-50"). It did what it should have done. I changed offsets to 34" and repeated. Warning on the left side - not unexpected, 50" not hard and fast limit. The images so far look fine - no jumps.

11:38 Resending, this time an active preset. Lots of warnings about copointing limits, but as expected since the offset is skirting the edge of the 50" copointing radius.

Longslit alignment (doing this on robs)

Note - I forgot to revert PCSInstrument.conf (the offset in RZP was 0.15 deg)

11:51 acqBinoMODS NGC40_Blue_Acq.txt NGC40_Red_Acq.txt

MODS2 - the thru-slit confirmatory image has no mask in. 55 should have been the confirmatory through slit image, but there is no slit in place. There was no error message and maskpos = IN.

Then I tried to put it in the focal plane and got an error:

2021-09-07T12:05:25.741] M2.IE> ERROR: SLITMASK SLITMASK=5 Impossible Mask in Science Field, but Grabber Stowed MASKPOS=IN GRABBER=STOW

41 deg elevation, rotator at 279 deg. We are stopping the tracking and rotator.

lbto@robs SpectroPhotSTDS ]$ isisCmd --mods2 m2.ie minsert rdbits
DONE: MINSERT SLITMASK=5 BITS=1001 b24-b21

10=mask is out

01 = grabber is in

Steve is slewing to zenith. The rotator is still at 279.

I reverted PCSInstrument.conf. The one I had created I wrote as CSInstrument.conf_new_mods1_rzp_20210907

Proposed Plan of Action:

  • MODS
    • MODS Rotator Center measure (use M71 at beg of night or M52 later)
      • requires pointing mask (40" square field stop & dot-circle with radius 45") - high (0.5hr)
    • MODS AGW 5/6 - Measure Rotator zero point (mask alignment) - high (0.5 hr) - mask exchange required
      • M1: dPA =
      • M2: dPA =
      • ssh -X into obs1 as telescope; then ssh - X as tcs@tcs1
        • The files needed are
          • /home/telescope/TCS/Configuration/GCS/MODS_{L,R}.cfg
          • /home/telescope/TCS/Configuration/PCS/PCSInstrument.conf
          • image files: /lbt/data/share/tcs/lbto.YYYMMDD
    • MODS AGW Transform data - ~1" seeing - high (2hours to collect data on-sky, time during day to analyze, ~0.5 hr-1hr to verify on-sky) - low priority
    • Spectrophotometric standard all modes - medium
      • Grating
      • Prism
    • non-sidereal guiding
      • Imaging
      • Long-slit - low priority
    • Test Science OB
      • longslit
      • imaging
        • Test without syncoffset and large dither (+/-45")
        • Test with synoffset and large dither (+/-45")
    • Dithering Repeatability with Pointing Mask to check DET XY
    • Skyflats
    • Blind offset tests
    • Pointing Model - 2hr
    • Test offsets for MODS1 for better alignment - requires test OT scripts (J. Power)
    • GCS and IDL have opposite field correction (IT 8315) - Medium/High (1-1.5hrs)
      • Requires 1" or better
      • run IDL task field_collect (same as measuring field aberrations)
      • take LUCI data first to cross-check IDL vs GCS
  • LBCs (if clear any seeing)
    • Standard photometric fields (new ones ready) - requires clear skies (any seeing) - required for ZP calibrations
    • Spectro-photometric standards - requires clear skies (any seeing) - required for ZP calibrations
    • Twilight Flats all filters - High (4 x 0.5 hrs - likely min two nights to complete)
    • Poor Conditions
      • LBC Pointing Model - (3hr) - any conditions
        • Standard Method
          • <a href='https://wiki.lbto.org/bin/edit/Commissioning/PtStarrs?topicparent=Commissioning.Com2021Restart' rel='nofollow' title='Create this topic'> PtStarrs </a> - please choose targets > 80 deg elevation (more than 1 or 2)
          • Run TMS passively - run 1 reference at the beginning
          • Need stable temperatures
          • Turn on PCS logging
          • Need to collimate at each star
          • Chip 2 only exposure
  • LUCI-1 (Medium)
    • Latest Scripts V1.31
    • Imaging Check
      • dither in field
        • (NGC 7027) - first half
        • - 2nd half (Scripts TBD)
      • non-sideral imaging dither in field (Scripts TBD)
    • Pointing Model -
      • PtStarrs ~10th mag 12-20 stars
      • Poor seeing/clouds OK
    • Photometric Standards (clear any seeing)


Current Status

  • LBCs
    • LBC-Blue - available
    • LBC-Red - available
  • MODS
    • MODS-1 - will not be at temp by night - possibly still OK for engineering?
      • AGW-5 - available
    • MODS-2 - available
      • AGW-6 - available
  • LUCI-1 - available
    • AGW-1 available
  • LUCI-2 offline
    • AGW-2 available
  • PEPSI - offline Software environmental system issues - earliest Monday
    • AGW-7/8 available


Closed Dome Activities

  • MODS-1 Full set of Cals
    • Biases
      • 1x1
      • 1x2
      • 2x2
    • Grating
      • Dual
      • Red
      • Blue
    • Imaging
      • Dual
      • Red
      • Blue
    • Prism
      • Dual
      • Red
      • Blue
  • MODS-2 Cals
    • Bias
      • 1x2
      • 2x2
    • Prism
      • Dual
      • Red
      • Blue
  • LUCI-1 Cals
    • Imaging N375 Flats
    • Narrow, medium-band filters
    • Broad zJHKKs and HeI completed
  • ARGOS Cal-Unit test for M3 alignment and pupil wobble checkout (alternate to on-sky method)
  • LUCI-1 Only Calibrations
    • Flexure data for G150
      • /home/dthomspons/Luci/FlexureCompensation
      • README - Calib15 - 2hrs
      • Ks filter, Argon lamp, optical sieve mask
      • take dark frame
      • Manual Control MCSGUI for elevations
    • Wavelength data for G150
      • G150 Ks, Argon, take image through slit, then put in grating, line should show up where the slit is located (within a few pixels of the image of the slit)
      • Check a few lines (2 or 3)

-- BarryRothberg - 06 Sep 2021
Topic revision: r9 - 16 Sep 2021, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback